OpenStudy (anonymous):

PLEASE HELP ME AND WILL GIVE MEDAL! Add the rational expressions 2x-1/(x-3)(x+2)+x-4/x-3

jagr2713 (jagr2713):

is there an answer choice

OpenStudy (anonymous):

common denominator will be \((x-3)(x+2)\)

OpenStudy (anonymous):

so \[\frac{2x-1}{(x-3)(x+2)}+\frac{x-4}{x-3}\] \[=\frac{2x-1+(x+2)(x-4)}{(x-3)(x+2)}\]

jagr2713 (jagr2713):

@satellite73 lol